# A Unified Framework for Diffusion Bridge Problems: Flow Matching and Schrödinger Matching into One

### v1

@@ 2.3 Iterative Proportional Filtering (IPF)
[…]
@@ 2.4 Iterative Markovian Fitting (IMF)
=$\mathcal{M}=\{P:dx_{t}=g_{t}(x_{t})dt+\sigma_{ref}dW_{t}\ \ \textrm{for any vector field}\ g\}$ (19)
@@ 3.1 A Special Case: (Mini-batch) Optimal Transport CFM (Tong et al, 2023)
where $\sigma_{min}\to 0$, and ($(\{x^{i}_{0}\}_{i\in B_{0}},\{x^{j}_{1}\}_{j\in B_{1}})$ is the mini-batch data, and $\{p^{mOT}_{ij}\}_{ij}$ is a $(|B_{0}|\times|B_{1}|)$ mini-batch OT solution matrix learned with the mini-batch as training data. That is,
$p^{mOT}=\arg\min_{p}\sum_{i,j}p_{ij}||x^{i}_{0}-x^{j}_{1}||^{2}\ \ \ \ \textrm{s.t.}\ \ \sum_{j\in B_{1}}p_{ij}=\frac{1}{|B_{0}|},\ \sum_{i\in B_{0}}p_{ij}=\frac{1}{|B_{1}|}$ (29)
It is easy to see that $P_{0}(\cdot)\to\pi_{0}(\cdot)$ and $P_{1}(\cdot)\to\pi_{1}(\cdot)$ as $\sigma_{min}\to 0$. Note also that $P_{t}(x|x_{0},x_{1})$ is always fixed over iterations while $Q(x_{0},x_{1})$ varies over iterations depending on the mini-batch data sampled. Note that in our UBA framework, each iteration allows for different choices of $P_{t}(x|x_{0},x_{1})$ and $Q(x_{0},x_{1})$.
@@ 3.2 A Special Case: (Mini-batch) Schrödinger Bridge CFM (Tong et al, 2023)
where ($(\{x^{i}_{0}\}_{i\in B_{0}},\{x^{j}_{1}\}_{j\in B_{1}})$ is the mini-batch data, and $\{p^{mEOT}_{ij}\}_{ij}$ is a $(|B_{0}|\times|B_{1}|)$ is the mini-batch entropic OT solution matrix with the negative entropy regularizing coefficient $2\sigma_{ref}^{2}$ (e.g., from the Sinkhorn-Knopp algorithm) learned with the mini-batch as training data. It is easy to see that the boundary conditions are met: $P_{0}(\cdot)=\pi_{0}(\cdot)$ and $P_{1}(\cdot)=\pi_{1}(\cdot)$.
@@ Appendix A Theorems and Proofs
    Practically, the training dynamics of the gradient descent for step 3 of Alg. 1 can be shown to be identical to that of minimizing $\mathbb{E}||v_{\theta}(t,x)-u_{t}(x)||^{2}$. This is done in Lemma A.3 although the proof is very similar to the result in Tong et al (2023). Hence, in practice, even without the assumptions of the ideal rich neural network functional capacity and perfect optimization, we can continue to reduce the error between $v_{\theta}(t,x)$ and $u_{t}(x)$ in the course of gradient descent for step 3.
@@
where its optimizer $v^{*}(t,x)$ equals the optimizer $v_{\theta^{*}}(t,x)$ of (51). In the functional optimization (58), the objective is completely decomposed over $t$, and we can equivalently minimize $\mathbb{E}_{Q(x_{0},x_{1})P_{t}(x_{t}|x_{0},x_{1})}||u_{t}(x_{t}|x_{0},x_{1})-v(t,x_{t})||^{2}$ for each $t$. We take the functional gradient with respect to $v(t,\cdot)$. For ease of exposition, we will use simpler notation where we minimize $\mathbb{E}_{p(y,z)}||f(y,z)-g(y)||^{2}$ with respect to the function $g(\cdot)$. Hence there is direct correspondence: $y$ is to $x_{t}$, $z$ to $(x_{0},x_{1})$, $f(y,z)$ to $u_{t}(x_{t}|x_{0},x_{1})$, and $g$ to $v$. We see that at optimum,
@@
$\nabla_{\theta}\mathbb{E}_{P_{t}(x)}||v_{\theta}(t,x)-u_{t}(x)||^{2}$ $=\nabla_{\theta}\mathbb{E}_{P_{t}(x)}\Big{[}||v_{\theta}(t,x)||^{2}-2v_{\theta}(t,x)^{\top}u_{t}(x)\Big{]}$ (60)
$=\nabla_{\theta}\mathbb{E}_{P_{t}(x)}||v_{\theta}(t,x)||^{2}-2\nabla_{\theta}\mathbb{E}_{P_{t}(x)}\Big{[}v_{\theta}(t,x)^{\top}u_{t}(x)\Big{]}$ (61)
@@
$=\nabla_{\theta}\int||v_{\theta}(t,x)||^{2}P_{t}(x|x_{0},x_{1})Q(x_{0},x_{1})d(x,x_{0},x_{1})$ (63)
$=\nabla_{\theta}\mathbb{E}_{P_{t}(x|x_{0},x_{1})Q(x_{0},x_{1})}||v_{\theta}(t,x)||^{2}$ (64)
The second term can be derived as:
@@
In step 2, since we set $u_{t}(x|x_{0},x_{1})=\alpha_{t}(x|x_{0},x_{1})$ where $\alpha$ is the solution to (3839), we will use the notation $u$ in place of $\alpha$ throughout the proof. First, we show that the SDE $dx_{t}=u_{t}(x_{t}|x_{0},x_{1})dt+\sigma dW_{t}$ with initial state $x_{0}$ at $t\!=\!0$ and $u$ satisfying (39), admits $\{P_{t}(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t};\mu_{t},\gamma_{t}^{2}I)\}_{t}$ as marginal distributions. Note that we must have $\mu_{0}=x_{0}$, $\mu_{1}=x_{1}$, $\gamma_{0}\to 0$, and $\gamma_{1}\to 0$ due to the conditioning (pinned process). This fact is in fact an extension of the similar one for ODE cases in (Tong et al, 2023). We will do the proof here for SDE cases. We will establish the Fokker-Planck equation for the SDE, and we derive:
@@
$=P_{t}(x|x_{0},x_{1})\cdot\Bigg{(}-\frac{\gamma_{t}^{\prime}}{\gamma_{t}}d+\frac{(x-\mu_{t})^{\top}\mu_{t}^{\prime}}{\gamma_{t}^{2}}+\frac{||x-\mu_{t}||^{2}\gamma_{t}^{\prime}}{\gamma_{t}^{3}}\Bigg{)}$ (70)
@@
$\ \ \ \ \ \ \ \ \ \ \ \ -P_{t}(x|x_{0},x_{1})\cdot\Bigg{(}-\frac{\gamma_{t}^{\prime}}{\gamma_{t}}d+\frac{(x-\mu_{t})^{\top}\mu_{t}^{\prime}}{\gamma_{t}^{2}}+\frac{||x-\mu_{t}||^{2}\Big{(}\gamma_{t}^{\prime}-\frac{\sigma^{2}}{2\gamma_{t}}\Big{)}}{\gamma_{t}^{3}}+\frac{\sigma^{2}}{2\gamma_{t}^{2}}d\Bigg{)}$ (71)
$\Delta P_{t}(x|x_{0},x_{1})=\textrm{Tr}\Big{(}\nabla_{x}^{2}P_{t}(x|x_{0},x_{1})\Big{)}=P_{t}(x|x_{0},x_{1})\cdot\Bigg{(}\frac{||x-\mu_{t}||^{2}}{\gamma_{t}^{4}}-\frac{d}{\gamma_{t}^{2}}\Bigg{)}$ (72)
@@
(Step 1 re-stated) Choose $P_{t}(x|x_{0},x_{1})$ as the marginals of the SDE, $dx_{t}=u_{t}(x_{t}|x_{0},x_{1})dt+\sigma dW_{t}$ with initial state $x_{0}$ at $t\!=\!0$ where $u$ is the solution to the constrained optimization:
@@
Let $P^{b}$ be the path measure of the SDE: $dx_{t}=b_{t}(x)dt+\sigma dW_{t}$, starting from the fixed initial state $x_{0}$. For the stochastic optimal control problem with the immediate cost $\frac{1}{2\sigma^{2}}||b_{t}(x_{t})||^{2}$ at time $t$ and the terminal cost $\log 1/g(x_{1})$ at final time $t=1$ for any function $g$, the cost-to-go function defined as:
$J_{t}^{b}(x):=\mathbb{E}_{P^{b}}\Bigg{[}\int_{t}^{1}\frac{1}{2\sigma^{2}}||b_{t}(x_{t})||^{2}-\log g(x_{1})\bigg{|}x_{t}=x\Bigg{]}$ (76)
@@
$\frac{\partial v_{t}(x)}{\partial t}+\frac{1}{2}\textrm{Tr}\Big{(}\sigma^{2}\nabla_{x}^{2}v_{t}(x)\Big{)}=\frac{\sigma^{2}}{2}||\nabla_{x}v_{t}(x)||^{2},\ \ v_{1}(\cdot)=-\log g(\cdot)$ (79)
has a solution $v_{t}(x)=-\log\mathbb{E}_{P^{ref}}[g(x_{1})|x_{t}=x]$. Note that we can write $\frac{\sigma^{2}}{2}||\nabla_{x}v_{t}(x)||^{2}$ as the following variational form,
$\frac{\sigma^{2}}{2}||\nabla_{x}v_{t}(x)||^{2}=-\min_{b}\ b^{\top}\nabla_{x}v_{t}(x)+\frac{||b||^{2}}{2\sigma^{2}}$ (80)
@@
$\frac{\partial v_{t}(x)}{\partial t}+\frac{1}{2}\textrm{Tr}\Big{(}\sigma^{2}\nabla_{x}^{2}v_{t}(x)\Big{)}=-\min_{b}\ b^{\top}\nabla_{x}v_{t}(x)+\frac{||b||^{2}}{2\sigma^{2}},\ \ v_{1}(\cdot)=-\log g(\cdot)$ (81)
Note that (81) is the Hamilton-Jacobi-Bellman equation for the stochastic optimal control problem with the immediate cost $\frac{1}{2\sigma^{2}}||b_{t}(x_{t})||^{2}$ and the terminal cost $\log 1/g(x_{1})$. In fact, $v_{t}(x)=-\log\mathbb{E}_{P^{ref}}[g(x_{1})|x_{t}=x]$ is the (optimal) value function, and the optimal control, i.e., the solution to (80) in a function form, is: $b^{*}_{t}(x)=-\sigma^{2}\nabla_{x}v_{t}(x)=\sigma^{2}\nabla_{x}\log\mathbb{E}_{P^{ref}}[g(x_{1})|x_{t}=x]$. ∎

### v2

@@ 2.3 Iterative Proportional Filtering (IPF)
$\textrm{(Reverse HB)}\ \ P^{2n-1}=\arg\min_{P}\ \textrm{KL}(P||P^{2n-2})\ \ \textrm{s.t.}\ P_{1}(x_{1})=\pi_{1}(x_{1})$ (15)
$\textrm{(Forward HB)}\ \ P^{2n}=\arg\min_{P}\ \textrm{KL}(P||P^{2n-1})\ \ \textrm{s.t.}\ P_{0}(x_{0})=\pi_{0}(x_{0})$ (16)
@@ 2.4 Iterative Markovian Fitting (IMF)
$\mathcal{M}=\{P:dx_{t}=g_{t}(x_{t})dt+\sigma_{ref}dW_{t}\ \ \textrm{for any vector field}\ g\}$ (19)
@@ 3.1 A Special Case: (Mini-batch) Optimal Transport CFM (Tong et al, 2023)
where $\sigma_{min}\to 0$, and $(\{x^{i}_{0}\}_{i\in B_{0}},\{x^{j}_{1}\}_{j\in B_{1}})$ is the mini-batch data, and $\{p^{mOT}_{ij}\}_{ij}$ is a $(|B_{0}|\times|B_{1}|)$ mini-batch OT solution matrix learned with the mini-batch as training data. That is,
[…]
It is worth mentioning that $\sigma_{min}\to 0$ is required to have boundary consistency for $P_{t}(x_{t}|x_{0},x_{1})$ at $t=0$ and $t=1$. Note also that $P_{t}(x|x_{0},x_{1})$ is always fixed over iterations while $Q(x_{0},x_{1})$ varies over iterations depending on the mini-batch data sampled. Note that in our UBA framework, each iteration allows for different choices of $P_{t}(x|x_{0},x_{1})$ and $Q(x_{0},x_{1})$.
@@ 3.2 A Special Case: (Mini-batch) Schrödinger Bridge CFM (Tong et al, 2023)
where $(\{x^{i}_{0}\}_{i\in B_{0}},\{x^{j}_{1}\}_{j\in B_{1}})$ is the mini-batch data, and $\{p^{mEOT}_{ij}\}_{ij}$ is a $(|B_{0}|\times|B_{1}|)$ is the mini-batch entropic OT solution matrix with the negative entropy regularizing coefficient $2\sigma_{ref}^{2}$ (e.g., from the Sinkhorn-Knopp algorithm) learned with the mini-batch as training data. Although the samples from the coupling distribution $Q(x_{0},x_{1})$ in (32) over the iterations in Alg. 1 conform to the data distributions $\pi_{0}$ and $\pi_{1}$ marginally, the mini-batch entropic OT solution is generally substantially different from the population entropic OT solution (the optimal solution of the Schrödinger Bridge).
@@ Appendix A Theorems and Proofs
[…]
@@
[…]
@@
The second term can be derived as:
$\nabla_{\theta}\mathbb{E}_{P_{t}(x)}\Big{[}v_{\theta}(t,x)^{\top}u_{t}(x)\Big{]}=\nabla_{\theta}\int v_{\theta}(t,x)^{\top}u_{t}(x)P_{t}(x)dx$ (65)
$\ \ \ \ \ \ =\nabla_{\theta}\int v_{\theta}(t,x)^{\top}\mathbb{E}_{Q(x_{0},x_{1})}[u_{t}(x|x_{0},x_{1})P_{t}(x|x_{0},x_{1})]dx$ (66)
@@
In step 2, since we set $u_{t}(x|x_{0},x_{1})=\alpha_{t}(x|x_{0},x_{1})$ where $\alpha$ is the solution to (3839), we will use the notation $u$ in place of $\alpha$ throughout the proof. First, we show that the SDE $dx_{t}=u_{t}(x_{t}|x_{0},x_{1})dt+\sigma dW_{t}$ with initial state $x_{0}$ at t=0t\!=\!0 and $u$ satisfying (39), admits $\{P_{t}(x_{t}|x_{0},x_{1})=\mathcal{N}(x_{t};\mu_{t},\gamma_{t}^{2}I)\}_{t}$ as marginal distributions. Note that we must have $\mu_{0}=x_{0}$, $\mu_{1}=x_{1}$, $\gamma_{0}\to 0$, and $\gamma_{1}\to 0$ due to the conditioning (pinned process). This fact is in fact an extension of the similar one for ODE cases in (Tong et al, 2023). We will do the proof here for SDE cases. We will establish the Fokker-Planck equation for the SDE, and we derive:
@@
[…]
@@
[…]
@@
(Step 1 re-stated) Choose $P_{t}(x|x_{0},x_{1})$ as the marginals of the SDE, $dx_{t}=u_{t}(x_{t}|x_{0},x_{1})dt+\sigma dW_{t}$ with initial state $x_{0}$ at t=0t\!=\!0 where $u$ is the solution to the constrained optimization:
@@
[…]
@@
[…]
@@
[…]